OpenStudy (goldphenoix):

If I multiply all whole numbers from 1 through 100, the largest power of 4 that is a factor of the product is A: 4^25 B: 4^32 C: 4^48 D: 4^50

terenzreignz (terenzreignz):

Where do you get these questions? -.-

OpenStudy (goldphenoix):

Lol. From a contest I took this year.

OpenStudy (goldphenoix):

Sadly, I failed the contest. Contest too hard. T~T

jimthompson5910 (jim_thompson5910):

hint: list all the multiples of 4 from 1 to 100

OpenStudy (goldphenoix):

LIST all them multiples of from 1 to 100? There's 25 number that is a multiple of 4.

terenzreignz (terenzreignz):

LOL Okay, annoying...

terenzreignz (terenzreignz):

I'm quite certain of my answer, but explaining it... gah...

terenzreignz (terenzreignz):

Say we have the product from 1 to 100 of all integers...

jimthompson5910 (jim_thompson5910):

you'll have 25 multiples of 4 you'll also have 50 multiples of 2....they pair up (in any way you want) to have 25 pairs which are multiples of 2*2 = 4 so you really have 25+25 = 50 copies of 4 being multiplied

OpenStudy (goldphenoix):

Yep.

terenzreignz (terenzreignz):

In fact, let's have it in prime factors...

terenzreignz (terenzreignz):

and now, let's take away all the 2's Actually, we need to count how many times 2 is multiplied in the factorisation of the product of all integers from 1 to 100.

OpenStudy (anonymous):

@jim_thompson5910 are you accounting for the fact that half the multiples of 2 are already accounted for as multiples of 4? And that a factor of 2 is only 4^(1/2)? So shouldn't we have: 25 multiples of 4 50 multiples of 2 - 25 multiples of 4 already counted = 25 new multiples of 2 = 12.5 multiples of 4?

jimthompson5910 (jim_thompson5910):

oh true

OpenStudy (anonymous):

If I absolutely had to guess I would say C. I'm honestly not sure how I would solve this problem.

jimthompson5910 (jim_thompson5910):

so (multiples of 2) + (multiples of 4) - (multiples of 2 and 4) = (25)+(25) - (25) = 25

terenzreignz (terenzreignz):

So... there are 50 multiples of 2, so that gives us 50 2's. There are 25 multiples of 4, so that gives us a FURTHER 25 2's There are 12 multiples of 8, so that gives us a FURTHER 12 2's There are 6 multiples of 16, so that gives us a FURTHER 6 2's There are 3 multiples of 32, so that gives us a FURTHER 3 2's There is 1 multiple of 64, so that gives us a FURTHER 1 2. Add those up, you get 50+25+12+6+3+2 = 96 2's So \[\Large 2^{96}=4^{48}\]is your man :D

OpenStudy (anonymous):

I think he nailed it

OpenStudy (goldphenoix):

Wowow. You too good TJ.

terenzreignz (terenzreignz):

It happens :D

OpenStudy (goldphenoix):

bow to TJ* I want to be your student. :3

terenzreignz (terenzreignz):

LOL that'll be the day :D

OpenStudy (goldphenoix):

:D